SARALANCE:

I need immediate assistance to the math problem 3x+2=14. Can anyone help?

TheKayla:

Let's solve your equation step-by-step. 3x+2=14 Step 1: Subtract 2 from both sides. 3x+2−2=14−2 3x=12 Step 2: Divide both sides by 3. 3x 3 = 12 3 x=4

SARALANCE:

@thekayla wrote:
Let's solve your equation step-by-step. 3x+2=14 Step 1: Subtract 2 from both sides. 3x+2−2=14−2 3x=12 Step 2: Divide both sides by 3. 3x 3 = 12 3 x=4
@TheKayla Thanks! 😄